OpenStudy (anonymous):

What is the square of 16? @iGreen. @iGreen.

jimthompson5910 (jim_thompson5910):

Hint: think of a square with a side length of 16  the area of this square is the same saying "the square of 16"

OpenStudy (anonymous):

4 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

if you wanted the square ROOT of 16, then yeah, \[\Large \sqrt{16} = 4\]

jimthompson5910 (jim_thompson5910):

but you asked about the square of 16 (I don't see "root" anywhere in your question)

OpenStudy (anonymous):

64

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the area of that square I drew above?

OpenStudy (anonymous):

16*16*16*16=

jimthompson5910 (jim_thompson5910):

you're thinking of the perimeter, which is 16+16+16+16 = 64

jimthompson5910 (jim_thompson5910):

I want the area

jimthompson5910 (jim_thompson5910):